# Time-Dependent Deflection Reconstruction: A New Technique to Search for Gravitational Waves with the Cosmic Microwave Background

### v1

@@ Abstract
Gravitational waves (GWs) passing through the Earth cause a correlated pattern of time-dependent deflections of the apparent position of astronomical sources. We build upon standard lensing reconstruction techniques to develop a new time-dependent quadratic estimator, providing a novel technique to search for the deflections produced by GWs using observations of the cosmic microwave background (CMB). We find that the time-dependent deflection reconstruction is many orders of magnitude more sensitive than the ordinary static lensing estimator, and it can be employed with the data collected by existing and future CMB surveys, without requiring any modification to the experimental or survey design. We demonstrate that CMB surveys offer sensitivity to GWs across a broad frequency range: while the sensitivity will not be competitive over the frequency range covered by pulsar timing arrays, it does extend coverage to both lower and higher frequencies. Finally, we discuss how our methods can be extended to search for other time-varying signals, and also how it can be applied to surveys at other wavelengths.
@@ II Impact of Gravitational Waves on the CMB
We start by considering the deflections produced by GWs and how they differ from gravitational lensing by large-scale structure in a few key aspects. First, the pattern of deflections due to GWs is expected to produce equal power in divergence-type ($E$-mode) and curl-type ($B$-mode) deflection [22], whereas gravitational lensing due to large-scale structure only produces divergence-type lensing (except on small angular scales where post-Born lensing corrections can become relevant [46, 47]). Next, the deflection due to GWs peaks on large angular scales and falls rapidly with multipole, reflecting the primarily quadrupolar nature of the deflection. This characteristic angular dependence of GW deflection is analogous to the Hellings-Downs curve, which describes the spatial correlation of pulsar timing residuals in the presence of GWs  [48]. Finally, and most essentially, the deflection due to a SGWB is time dependent, while the lensing by large-scale structure remains static over the duration of the survey, changing only on cosmological time scales (and similarly the primary CMB anisotropies evolve very slowly compared to the survey duration [49, 50, 51]). Figure 1 shows the effect of GWs on a simulated map of the CMB.
@@
where $H_{0}$ is the Hubble constant and $\Omega_{\mathrm{GW}}(f)$ is the energy density in GWs per unit logarithmic frequency divided by the critical energy density, i.e. $\bar{\Omega}_{\mathrm{GW}}=\int d\log f\,\Omega_{\mathrm{GW}}(f)$, and we use an overbar to denote quantities integrated over frequency. Furthermore, the deflections produced by these GWs can be broken down into harmonic modes $L$, divergence- and curl-type deflections $\Gamma\in[\phi,\psi]$, and GW temporal frequencies $f$ according to
$\theta_{\mathrm{rms}}^{2}(f)=\sum_{L=2}^{\infty}\sum_{\Gamma=\phi,\psi}\frac{g_{\Gamma}\,\alpha_{L}^{\Gamma\Gamma}\,H_{0}^{2}\,\Omega_{\mathrm{GW}}(f)}{4\pi^{2}\,f^{2}}\,$, (2)
@@
$\langle\mathcal{T}^{N}(\mathbf{l}_{1},\omega_{1})\mathcal{T}^{N}(\mathbf{l}_{2},\omega_{2})^{\star}\rangle$
@@
$N_{l}^{TT}$ $=\Delta_{T}^{2}\exp\left(l(l+1)\frac{\theta_{\mathrm{fwhm}}}{8\log 2}\right)\,$, (5)
@@ III Time-Dependent Deflection Reconstruction
$A^{\Gamma}(L,\Omega)=\left[\int\frac{d^{2}\mathbf{l}_{1}}{(2\pi)^{2}}\mathbf{l}_{1}\odot\mathbf{L}\,\tilde{\mathcal{C}}^{\mathcal{TT}}_{l_{1},0}F^{\Gamma}(\mathbf{l}_{1},\mathbf{l}_{2},\Omega)\right]^{-1}\,$, (14)
and we have introduced the operator $\odot$ to represent a dot product when $\Gamma=\phi$ and a dot product with the ninety-degree rotated vector when $\Gamma=\psi$ [60].
@@
$F^{\Gamma}(\mathbf{l}_{1},\mathbf{l}_{2},\Omega)$ $=\frac{\mathbf{l}_{1}\odot(\mathbf{l}_{1}+\mathbf{l}_{2})\,\tilde{\mathcal{C}}^{\mathcal{TT}}_{l_{1},0}}{\mathcal{C}_{l_{1},0}^{\mathcal{TT},\mathrm{obs}}\mathcal{C}_{l_{2},\Omega}^{\mathcal{TT},\mathrm{obs}}}=\frac{f^{\Gamma}(\mathbf{l}_{1},\mathbf{l}_{2},0,\Omega)}{\mathcal{C}_{l_{1},0}^{\mathcal{TT},\mathrm{obs}}\mathcal{C}_{l_{2},\Omega}^{\mathcal{TT},\mathrm{obs}}}\,$, (16)
where the second equality comes from the fact that the undeflected temperature is time independent $\tilde{\mathcal{C}}_{l,\omega}^{\mathcal{TT}}=2\pi\delta(\omega)\tilde{C}_{l}^{TT}$. The 3d noise power of our time-dependent deflection reconstruction is then
@@
where in the last equality, we assumed $\Omega\neq 0$ (and again used that $2\pi\delta(0)=\Delta t_{\mathrm{survey}}$). Note that if we had taken $\Omega=0$, we would recover the ordinary static lensing estimator [40]. In the case of white noise in temperature, we find $\mathcal{N}_{L,\Omega}^{\phi\phi}=\mathcal{N}_{L,\Omega}^{\psi\psi}$. The estimators for $\phi$ and for $\psi$ are nearly orthogonal, though there is a very small cross talk between the two estimators from the parts of the temperature spectrum where the scale dependence deviates from power-law behavior [60]; we will neglect this small departure from orthogonality in what follows.
@@
In Figure 3, we show the time-dependent deflection reconstruction 3d noise power in a frequency band of width $\Delta t_{\mathrm{survey}}^{-1}$ and the static deflection reconstruction 2d noise power $N_{L}^{dd}$ (which is the same as the time-dependent deflection 3d noise power in a frequency band of width $\Delta t_{\mathrm{survey}}^{-1}$ at $\Omega=0$, $N_{L}^{dd}=\Delta t_{\mathrm{survey}}^{-1}\mathcal{N}_{L,0}^{dd}$), and we have defined the deflection noise power in terms of the deflection potential noise power as $\mathcal{N}_{L,\Omega}^{dd}=L^{2}\mathcal{N}_{L,\Omega}^{\phi\phi}=L^{2}\mathcal{N}_{L,\Omega}^{\psi\psi}$. We obtain these by evaluating Eq. (17), using the symlens code [62] with our custom estimator. Since the noise power is independent of $\Omega$—i.e., since the temporal white noise in the CMB leads to white noise in the reconstructed deflections—we implement the 2d version of the integrals, namely, the last line of Eq. (17).
@@
We could formulate our time-dependent deflection reconstruction using different polarization combinations as well. However, we focused here on the $TT$ estimator for time-dependent deflection since it has better performance than estimators constructed from $E$ and $B$. This is due to the fact that the time-dependent estimator benefits from having as many high signal-to-noise modes as possible on the $\omega=0$ leg, and as low a noise as possible on the $\Omega\neq 0$ leg. The time-independent signal for $T$ is significantly larger than $E$ or $B$ (from both primary CMB fluctuations and from astrophysical foregrounds), and the noise is also lower in $T$ for the small-scale CMB observations we consider here. Therefore, unlike for static lensing reconstruction, where the $EB$ estimator is expected to provide the best performance at low noise (owing to a similar lack of cosmic variance in the unlensed $B$ field) [40], the $TT$ estimator exhibits the best performance for time-dependent deflection reconstruction.
@@ V.2 Integrated constraint
The deflections caused by GWs with periods longer than the duration of the survey will cause a secular drift in astrophysical source positions. The contribution to the mean square angular velocity due to a GW of frequency $f$ is related to the mean square deflection by $\dot{\theta}_{\mathrm{rms}}^{2}(f)=f^{2}\theta_{\mathrm{rms}}^{2}(f)$, and the observed angular velocity receives contributions from all frequencies [19, 22]. The mean square angular velocity of source positions due to GWs below the resolvable band is
$\left(\dot{\theta}_{\mathrm{rms}}^{\mathrm{low}}\right)^{2}=\frac{1}{4\pi^{2}}\int_{f<\Delta t_{\mathrm{survey}}^{-1}}d\log f\,H_{0}^{2}\Omega_{\mathrm{GW}}(f)\,$, (33)
@@
$L^{3}f^{3}\mathcal{C}_{L,(2\pi f)}^{\Gamma\Gamma}=g_{\Gamma}\alpha_{L}^{\Gamma\Gamma}\left(\dot{\theta}_{\mathrm{rms}}^{\mathrm{low}}\right)^{2}\,.$ (34)
@@ Acknowledgments
The authors thank Colin Hill and Mathew Madhavacheril for helpful discussions. Computational resources for this research were provided by SMU’s O’Donnell Data Science and Research Computing Institute. JM and AvE are supported by NASA through ADAP grant 80NSSC24K0665. JM is additionally supported by the US Department of Energy under Grant DE-SC0010129. AvE is additionally supported by NASA ADAP grants 80NSSC23K0747 and 80NSSC23K0464, and NSF AAG grant 588167.

### v2

# Time-dependent deflection reconstruction: new technique to search for gravitational waves with the cosmic microwave background

@@ Abstract
Gravitational waves (GWs) passing through the Earth cause a correlated pattern of time-dependent deflections of the apparent position of astronomical sources. We build upon standard lensing reconstruction techniques to develop a new time-dependent optimal quadratic estimator, providing a novel technique to search for the deflections produced by GWs using observations of the cosmic microwave background (CMB). We find that the time-dependent deflection reconstruction is many orders of magnitude more sensitive than the ordinary static lensing estimator, and that it can be employed with the data collected by existing and future CMB surveys, without requiring any modification to the experimental or survey design. We demonstrate that CMB surveys offer sensitivity to GWs across a broad frequency range: while the sensitivity will not be competitive over the frequency range covered by pulsar timing arrays, the coverage band extend to both lower and higher frequencies. Finally, we discuss how our methods can be adapted to search for other time-varying signals, and also how they can be applied to surveys at other wavelengths.
@@ II Impact of Gravitational Waves on the CMB
We start by considering the deflections produced by GWs and how they differ from gravitational lensing by large-scale structure in a few key aspects. First, the pattern of deflections due to GWs is expected to produce equal power in divergence-type ($E$-mode) and curl-type ($B$-mode) deflection [20], whereas gravitational lensing due to large-scale structure only produces divergence-type lensing (except on small angular scales where post-Born lensing corrections can become relevant [79, 65, 84]). Next, the deflection due to GWs peaks on large angular scales and falls rapidly with multipole, reflecting the primarily quadrupolar nature of the deflection. This characteristic angular dependence of GW deflection is analogous to the Hellings-Downs curve, which describes the spatial correlation of pulsar timing residuals in the presence of GWs  [46]. Finally, and most essentially, the deflection due to a SGWB is time dependent, while the lensing by large-scale structure remains static over the duration of the survey, changing only on cosmological time scales (and similarly the primary CMB anisotropies evolve very slowly compared to the survey duration [54, 99, 73]). Figure 1 shows the effect of GWs on a simulated map of the CMB.
@@
where $H_{0}$ is the Hubble constant and $\Omega_{\mathrm{GW}}(f)$ is the energy density in GWs per unit logarithmic frequency divided by the critical energy density, i.e. $\bar{\Omega}_{\mathrm{GW}}=\int d\log f\,\Omega_{\mathrm{GW}}(f)$, and we use an overbar to denote quantities integrated over frequency (such that $\bar{\theta}_{\mathrm{rms}}^{2}=\int d\log f\,\theta_{\mathrm{rms}}^{2}(f)$). The derivation of Eq. (1) presented in Ref. [20] proceeds by first computing the apparent deflection of a photon source due to an arbitrary GW signal $h_{ij}$, assuming that both the source and observer a freely falling in an expanding universe. Under the assumption that the photon source is at a distance much larger than the wavelength of the GW and that the GW wavelength is short compared to the horizon size, the deflection takes a simple form for plane GWs, depending only on the GW evaluated at the position of the observer
$\delta n^{i}(\tau,\hat{\mathbf{n}})=\frac{n^{i}+p^{i}}{2(1+\mathbf{p}\cdot\hat{\mathbf{n}})}h_{jk}(\tau,0)n_{j}n_{k}-\frac{1}{2}h_{ij}(\tau,0)n_{j}\,$, (2)
where $\hat{\mathbf{n}}$ denotes the direction in the sky, $\tau$ is the time of observation, the GW propagates along direction $\mathbf{p}$, and $h_{ij}(\tau,0)$ refers to the GW at the position of the observer. The apparent deflection of the source accounts for both the perturbed path of the photon and the perturbation to the observer’s local proper reference frame in the presence of the GW. The statistical properties of the deflection due to a SGWB can then be computed from the two-point correlation function of the metric perturbations, which can be treated as a superposition of plane waves.
The deflections produced by these GWs can be broken down into harmonic modes $L$, divergence- and curl-type deflections $\Gamma\in[\phi,\psi]$, and GW temporal frequencies $f$ according to
$\theta_{\mathrm{rms}}^{2}(f)$ $=\sum_{L=2}^{\infty}\sum_{\Gamma=\phi,\psi}\theta_{L,\Gamma}^{2}(f)$
$=\sum_{L=2}^{\infty}\sum_{\Gamma=\phi,\psi}\frac{g_{\Gamma}\,\alpha_{L}^{\Gamma\Gamma}\,H_{0}^{2}\,\Omega_{\mathrm{GW}}(f)}{4\pi^{2}\,f^{2}}\,$, (3)
@@
$\langle\mathcal{T}^{\mathrm{N}}(\mathbf{l}_{1},\omega_{1})\mathcal{T}^{\mathrm{N}}(\mathbf{l}_{2},\omega_{2})^{\star}\rangle$
@@
$N_{l}^{TT}$ $=\Delta_{T}^{2}\exp\left(l(l+1)\frac{\theta^{2}_{\mathrm{fwhm}}}{8\log 2}\right)\,$, (6)
@@ III Time-Dependent Deflection Reconstruction
$A^{\Gamma}(L,\Omega)=\left[\int\frac{d^{2}\mathbf{l}_{1}}{(2\pi)^{2}}f^{\Gamma}(\mathbf{l}_{1},\mathbf{l}_{2},0,\Omega)F^{\Gamma}(\mathbf{l}_{1},\mathbf{l}_{2},\Omega)\right]^{-1}\,.$ (15)
@@
$F^{\Gamma}(\mathbf{l}_{1},\mathbf{l}_{2},\Omega)$ $=\frac{f^{\Gamma}(\mathbf{l}_{1},\mathbf{l}_{2},0,\Omega)}{\mathcal{C}_{l_{1},0}^{\mathcal{TT},\mathrm{obs}}\mathcal{C}_{l_{2},\Omega}^{\mathcal{TT},\mathrm{obs}}}\,.$ (17)
The 3d noise power of our time-dependent deflection reconstruction is then
@@
where we assumed $\Omega\neq 0$ (and again used that $2\pi\delta(0)=\Delta t_{\mathrm{survey}}$). We have introduced the operator $\odot$ to represent a dot product when $\Gamma=\phi$ and a dot product with the ninety-degree rotated vector when $\Gamma=\psi$ [30]. Note that if we had taken $\Omega=0$, we would recover the ordinary static lensing estimator [48]. There is an important difference in the mode coupling that arises in the case of the time-dependent deflection and that of static deflection. In the limit where we utilize small-scale temperature modes to reconstruct large scale deflection potentials ($l_{1}\gg L$) such that $\tilde{\mathcal{C}}_{l_{1},0}^{\mathcal{TT}}\simeq\tilde{\mathcal{C}}_{l_{2},0}^{\mathcal{TT}}$, the mode coupling in the static case is proportional to the large-scale convergence $f^{\Gamma}(\mathbf{l}_{1},\mathbf{L}-\mathbf{l}_{1},0,0)\simeq L^{2}\tilde{\mathcal{C}}_{l_{1},0}^{\mathcal{TT}}$ such that $\langle\mathcal{T}\mathcal{T}\rangle\propto L^{2}\Gamma(L)$, whereas in the time-dependent case the mode coupling is proportional to the large-scale deflection $f^{\Gamma}(\mathbf{l}_{1},\mathbf{L}-\mathbf{l}_{1},0,\Omega\neq 0)\simeq(\mathbf{L}\odot\mathbf{l}_{1})\tilde{\mathcal{C}}_{l_{1},0}^{\mathcal{TT}}$ such that $\langle\mathcal{T}\mathcal{T}\rangle\propto L\Gamma(L)$. As a result, the reconstruction noise power is flat in convergence units for the static reconstruction, and flat in deflection units for the time-dependent reconstruction case. This distinction could be anticipated physically by noting that a static and spatially constant deflection is unobservable, though a time-dependent deflection can be observed, even if it is spatially constant [^1] We thank the anonymous referee for pointing out this important fact.. The different behavior of the reconstruction noise power on large scales is apparent below in Figure 3. In the case of white noise in temperature, we find $\mathcal{N}_{L,\Omega}^{\phi\phi}=\mathcal{N}_{L,\Omega}^{\psi\psi}$. The estimators for $\phi$ and for $\psi$ are orthogonal since the filters for the two estimators have opposite parity [75].
@@
In Figure 3, we show the time-dependent deflection reconstruction 3d noise power in a frequency band of width $\Delta t_{\mathrm{survey}}^{-1}$ and the static deflection reconstruction 2d noise power $N_{L}^{dd}$ (which is the same as the time-dependent deflection 3d noise power in a frequency band of width $\Delta t_{\mathrm{survey}}^{-1}$ at $\Omega=0$, $N_{L}^{dd}=\Delta t_{\mathrm{survey}}^{-1}\mathcal{N}_{L,0}^{dd}$), and we have defined the deflection noise power in terms of the deflection potential noise power as $\mathcal{N}_{L,\Omega}^{dd}=L^{2}\mathcal{N}_{L,\Omega}^{\phi\phi}=L^{2}\mathcal{N}_{L,\Omega}^{\psi\psi}$. We obtain these by evaluating Eq. (18), using the symlens code [60] with our custom estimator. (This expression is obtained in the flat sky formalism, but we extrapolate down to $L=2$ using the fact that the time-dependent deflection noise is independent of $L$ on large scales.) Since the noise power is independent of $\Omega$—i.e., since the temporal white noise in the CMB leads to white noise in the reconstructed deflections—we implement the 2d version of the integrals, namely, the last line of Eq. (18).
@@
We could formulate our time-dependent deflection reconstruction using different polarization combinations as well. However, we focused here on the $TT$ estimator for time-dependent deflection since it has better performance than estimators constructed from $E$ and $B$. This is due to the fact that the time-dependent estimator benefits from having as many high signal-to-noise modes as possible on the $\omega=0$ leg, and as low a noise as possible on the $\Omega\neq 0$ leg. The time-independent signal for $T$ is significantly larger than $E$ or $B$ (from both primary CMB fluctuations and from astrophysical foregrounds), and the noise is also lower in $T$ for the small-scale CMB observations we consider here. Furthermore, due to the angular dependence of the $EB$ mode coupling, the time-dependent deflection noise from the $EB$ reconstruction is flat in convergence units rather than deflection units on large angular scales, just as in the static case. Therefore, unlike for static lensing reconstruction, where the $EB$ estimator is expected to provide the best performance at low noise (owing to a similar lack of cosmic variance in the unlensed $B$ field) [48], the $TT$ estimator exhibits the best performance for time-dependent deflection reconstruction.
@@ V.2 Integrated constraint
The deflections caused by GWs with periods longer than the duration of the survey will cause a secular drift in astrophysical source positions. The contribution to the mean square angular velocity due to a GW of frequency $f$ is related to the mean square deflection by $\dot{\theta}_{\mathrm{rms}}^{2}(f)=4\pi^{2}f^{2}\theta_{\mathrm{rms}}^{2}(f)$, and the observed angular velocity receives contributions from all frequencies [81, 20]. The mean square angular velocity of source positions due to GWs below the resolvable band is
$\left(\dot{\theta}_{\mathrm{rms}}^{\mathrm{low}}\right)^{2}=\int_{f<\Delta t_{\mathrm{survey}}^{-1}}d\log f\,H_{0}^{2}\Omega_{\mathrm{GW}}(f)\,$, (34)
@@
$4\pi^{2}L^{3}f^{3}\mathcal{C}_{L,(2\pi f)}^{\Gamma\Gamma}=g_{\Gamma}\alpha_{L}^{\Gamma\Gamma}\left(\dot{\theta}_{\mathrm{rms}}^{\mathrm{low}}\right)^{2}\,.$ (35)
@@ Acknowledgments
The authors thank Colin Hill, Mathew Madhavacheril, and Toshiya Namikawa for helpful discussions. We also thank the anonymous referee who suggested a physical justification for the shape of the reconstruction noise curves in Figure 3. Computational resources for this research were provided by SMU’s O’Donnell Data Science and Research Computing Institute. JM and AvE are supported by NASA through ADAP grant 80NSSC24K0665 and by NSF through grant AST-2510926. JM is additionally supported by the US Department of Energy under Grant DE-SC0010129. AvE is additionally supported by NASA ADAP grants 80NSSC23K0747 and 80NSSC23K0464, and NSF AAG grant 588167.